# Digital currency hardware wallets and the essence of money

### v3

@@ 6 Wallets as containers for bearer instruments
For a set of specific technical requirements for digital currency in general to be credible, it must address the requirements of its various stakeholders. One of the unfortunate characteristics of the current debate is the diversity of conflicting assumptions about the nature of payments and the relative importance of different kinds of payment scenarios. In reality, there are many different ways that people make payments today: with and without custodians, locally and at a distance, with and without electronic devices, with and without money issued by a central bank. Money is transferred to pay invoices, to purchase goods and services, to give gifts, to pay taxes, or to reallocate funds within an organisation or family. These different scenarios imply a plethora of different costs and risk profiles: Buying a cup of coffee is different from buying a house. We should accept that there will not be a single payment system or key mechanism to rule them all, although it might be possible to develop a common framework that can accommodate a variety of different payment systems.
@@
The case for trusted hardware, as for other core mechanisms of proposed payment systems, involves the requirement for assessing the integrity of money in a fair exchange scenario wherein counterparties lack access to a mutually trusted third party. In principle, trusted hardware can act as a trusted third party to enforce system-level rules on user devices directly, by implementing rules that prevent assets from being spent more than once. However, considering the security risks associated with compromised devices, it is better to enforce such rules over the network instead, if it is possible to do so. This can be done, for example, by invoking a trusted third party to validate all transactions, as proposed by Chaum and others [33, 34]. However, there is reason to believe the costs, risks, and public accountability concerns associated with having a central operator dictate that history unchecked are too great for the CBDC use case [35], and the most important costs and risks would be borne not only by the users but also by the central operator. A modern alternative involves using a distributed ledger system to put the money itself on a blockchain, the way that UTXO designs (e.g. Bitcoin [32], Monero [39]) and state-transition designs (e.g. Ethereum [40]) do. Such approaches are unwieldy because they require the ledger to track all of the tokens individually, and transaction finality requires global consensus. In practice, most cryptocurrency users rely upon custodial accounts, negating a foundational benefit of digital currency.
However, distributed ledgers and blockchains can be used to support and enable payment systems in a variety of ways, including by providing non-repudiable evidence that a transaction had taken place, even if an individual service provider might equivocate [41]. The digital currency design proposed by Goodell et al [10] and further elaborated by Friolo et al [38] addresses the scalability problems associated with modern cryptocurrency systems while strongly preventing double spending without trusted hardware. In particular, it avoids requiring the issuer to participate in every transaction of its tokens, it avoids requiring system operators to manage a list of tokens that have been created or spent, and it avoids requiring users to have access to a global network at the time of the transaction. Any third party chosen by the spender and accessible to the recipient at the time of the transaction can suffice.
Proposed solutions that rely upon stateful assets and oblivious transfers can work technically without distributed ledgers, but in the end, digital currency systems require a means to support an immutable, shared history mutually accepted among independent actors, which is the salient characteristic of a distributed ledger [42]. The immutability of a distributed ledger is emergent property of the independence of the participants. So, although using a distributed ledger to record and process tokens directly on the ledger might bring unacceptable costs in terms of scale, efficiency, and operational burden, we should not rule out using distributed ledgers for what they do best: externalising commitments so that parties responsible for maintaining the history do not equivocate [10]. These commitments can be aggregated in such a way that the ledger grows at a fixed rate determined in advance, and each transaction carries zero marginal cost to the ledger operators.
@@
Trusted computing is not necessary to protect users from loss or theft, either. To protect their assets, users may, at their option, choose devices that offer security features such as encryption, passcodes, or local biometrics to protect against theft, and they may make copies of their tokens to mitigate the risk of accidental loss of their devices. Users can retain copies of their tokens, and they can share copies of the tokens with the police if the devices containing them are lost, stolen, or compromised. Users can also upload copies of their tokens to non-custodial service providers, such as cloud storage services.
@@
A much better way to address the limited set of cases in which “fully offline” payments (that is, those where neither the payer nor the payee have access to any suitable third parties or network infrastructure) are necessary would be to support the use of cash itself where cash continues to work well, such as in remote locations and environments with low connectivity, while avoiding all of the aforementioned problems associated with trusted hardware. To address the reduction of cash in the “built” environments by providing digital currency, whilst encouraging the provision of facilities for converting between cash and diital currency at the boundary of “unbuilt” environments, would be an entirely legitimate approach. In particular, if one potential benefit of CBDC in an age of deteriorating cash infrastructure is the ability to electronically transport cash to remote population centres during crises [43], then there is no reason to require that central bank money distributed in this manner must remain digital. For example, it can be exchanged for physical tokens specific to that population centre, much as cash and bank money are exchanged for chips at a casino.
@@
None of this is to say that we should forbid limited groups of users from trusting certified devices, if that is what they freely choose, to conduct transactions without network connectivity to a third party. Such an approach is entirely compatible with an architecture that expects that the third party would generally operate as a network service. However, at a system level, requiring all users to purchase and carry certified devices, even if some users intend only to hold money offline so that they can transact it online, introduces costs and risks that are entirely avoidable.
@@ 7 Conclusion
In the end, we must ask ourselves: What problem are we really trying to solve with digital currency? Manifestly, the problem is the decline of cash, and the concomitant challenges of maintaining cash infrastructure when the variable revenues of cash infrastructure fall below the fixed costs. But the decline of cash really only exists in environments with high network connectivity, wherein dual-offline payments are not necessary. For environments with low network connectivity, the decline of cash has not occurred, and cash continues to be the best solution for such environments.
@@
Many central banks, including the Bank of England, have affirmed a commitment to cash, while underscoring the risks associated with the decline of cash use in favour of online payment schemes such as card payment networks. Really, the use cases that digital currency payments should be targeting are cases for which at least one counterparty is already online. To argue for digital currencies with the premise that support for dual-offline payments are its principal source of added value mainly supports the view that such products are solutions in search of a problem.
@@
The purpose of digital wallets (and indeed, digital currency in general) is not primarily about enabling transactions in which both parties are offline. The purpose of digital wallets is to allow users to hold assets directly, without a custodian. Put differently, this is a way for users to have private property that they can exchange. By facilitating such payments, digital currency can protect the interests of consumers in all cases, including fully networked ones. If the counterparty has network connectivity (for example, an electronic point of sale device), then transactions involving digital wallets can be consummated without special hardware. Even if we accept that fair exchange is impossible without the participation of a trusted third party, and that the trusted third party can participate via the network or via trusted hardware, it is important to underscore that a non-networked digital wallet can conduct a fair exchange with a networked counterparty, without the aid of special security mechanisms built into the wallet.
@@
It is also a mistake to assume the primary motivation for CBDC is to allow a government to phase out cash, or that custodial wallets are sufficient for all scenarios except the narrow case in which neither the sender nor the receiver have a communication line to a third party. Part of the reason that cash is successful is its role as a bearer instrument. Credit systems actually predated cash, and cash emerged as a partial substitute [44]. What the world does not yet have is an institutionally supported version of cash for the digital economy, and a digital currency aspiring to fill this role will require a non-custodial wallet architecture.
@@
Custodial wallets require network access as a rule, because consumers must be able to use their devices to access their custodial accounts. However, digital wallets do not imply a design choice between network access and certified hardware in the general case. In particular, single-offline payments can make use of non-custodial wallets without certified hardware to hold money offline and transact it online, leveraging the network connectivity provided by POS devices presented to customers by merchants. To deny the salience of this use of wallets is to impede the development of a technology that may one day become crucial for the digital economy. Ultimately, the digital wallets that support economic transactions should be seen not as access devices or proxies offered by custodians, but as tools freely chosen, and even built, by end-users to store and use digital money autonomously, on their own terms.
@@ Acknowledgements
The author thanks Professor Tomaso Aste for his continued support for our work on digital currencies and digital payment systems. The author also thanks his many colleagues in the international standards development community for a series of vigorous arguments that were instrumental to reaching the crux of this most contentious debate. The author also acknowledge the Future of Money Initiative at University College London, the Systemic Risk Centre at the London School of Economics, EPSRC, and the PETRAS Research Centre EP/S035362/1 for the FIRE Project.

### v4

@@ 6 Wallets as containers for bearer instruments
possession self-custody versus third party custody
ID requirement yes (account) versus no (bearer)
proximity in-person (POS) versus remote (e-commerce)
device authorisation certified hardware versus public protocol
form physical (cash) versus digital
issuer central bank versus private sector entity
regulation bank money versus e-money
Table 3: Some dimensions by which payment technologies can be described.
@@
For a set of specific technical requirements for digital currency in general to be credible, it must address the requirements of its various stakeholders. One of the unfortunate characteristics of the current debate is the diversity of conflicting assumptions about the nature of payments and the relative importance of different kinds of payment scenarios. In reality, there are many different dimensions that can be used to describe how people make payments today: with and without custodians, locally and at a distance, with and without electronic devices, with and without money issued by a central bank; see Table 3. Money is transferred with finality333as a way of discharging obligations between transacting parties. for a variety of reasons, including to pay invoices, to purchase goods and services, to give donations, to pay taxes, to repay loans, or to transfer funds within an organisation or family. These different scenarios imply a plethora of different costs and risk profiles: Buying a cup of coffee is different from buying a house. We should accept that there will not be a single payment system or key mechanism to rule them all, although it might be possible to develop a common framework that can accommodate a variety of different payment systems.
@@
The case for trusted hardware, as for other core mechanisms of proposed payment systems, involves the requirement to ensure the integrity of money when counterparties lack access to a mutually trusted third party. In principle, trusted hardware can act as a trusted third party to enforce system-level rules on user devices directly, by implementing rules that prevent digital assets held in a wallet from being spent more than once. However, considering the security risks associated with compromised devices, it is better to enforce such rules over the network instead, if it is possible to do so. This can be done, for example, by requiring the issuer to validate all transactions, as proposed by Chaum and others [Chaum 1982, Chaum 2021]. However, there is reason to believe the costs, risks, and public accountability concerns associated with having a single central operator perform issuance, settlement, and compliance functions might be too great for the CBDC use case [Goodell 2021].
A modern alternative to the centralised model involves using a distributed ledger system to put the money itself on a blockchain, the way that UTXO designs (e.g. Bitcoin, Monero [Van Saberhagen 2013]) and state-transition designs (e.g. Ethereum [Buterin 2013]) do. However, such approaches are unwieldy because they require the ledger to track all of the tokens individually, and transaction finality requires global consensus, introducing a system-level economic externality with every transaction, exposing users to significant management costs, and, in some cases, creating uncertainty about finality. In practice, most cryptocurrency users rely upon custodial accounts, negating a foundational benefit of digital currency.
feature UTXO (e.g. Bitcoin) Chaum (GNU Taler) Oblivious management
assets represented directly on a ledger when issued yes no no
assets represented directly on a ledger when spent yes yes no
bundled functions shared across network peers yes no no
finality assurance supported by network peers yes no optional
issuance separate from settlement and compliance no no yes
settlement separate from compliance no no yes
Table 4: Features of selected token-based payment systems.
And so, instead of aiming to share the responsibility for a bundled set of issuance, settlement, and compliance enforcement functions across a set of peers, it might be better to decompose the functions into distinct roles that can be performed by different parties, enabling the costs and risks to be allocated more efficiently. Decentralised infrastructure can be used to support and enable payment systems in a variety of ways, including by providing non-repudiable evidence that a transaction had taken place, even if an individual service provider might equivocate [Goodell 2026]. Table 6 provides a comparative overview of the various token-based payment systems that we have mentioned in terms of their key features.
The digital currency design proposed by Goodell et al [Goodell 2023] and further elaborated by Friolo et al [Friolo 2025] offers an alternative approach to achieving scalability, replacing global consensus about the status and ownership of individual assets, and the related costs of such consensus common to modern cryptocurrency systems, with a combination of local trust and a mechanism for keeping local trust providers honest. By integrating a fungible token created using Chaum’s method that delivers privacy by design for consumers into a non-fungible asset that can be transferred with strong traceability following the initial creation, it enables the separation of the costs and risks associated with the aforementioned roles that enable payments. With this design, the third parties that provide transaction finality do not manage tokens, the issuer is not involved in individual transactions, and regulated entities, such as banks, can continue to provide compliance enforcement (e.g. income monitoring for the purposes of taxation and sanctions enforcement) as a condition of exchanging a spent token for a fresh token or another form of money.
Oblivious management can work technically with a centralised notary, but in the end, digital currency systems require common agreement about history, which is the salient characteristic of a distributed ledger [ISO 22739 2024]. Agreement about history does not imply shared knowledge of the holdings or transfers of individual assets. If assets maintain their own provenance information with references to ledger state, then having an integrity provider record only the zero-knowledge commitments of payers to their transactions, and having the ledger record only the rolled-up aggregation of such commitments, is sufficient. These commitments can be aggregated in such a way that the ledger grows at a fixed rate determined in advance, without individual transactions introducing system-level externalities, and such that each transaction carries zero marginal cost to the ledger operators. Some established DLT systems, such as Arbitrum, already aggregate transactions in this way [Kalodner 2018]. Proof that a transaction is part of a fixed-size aggregation, which, subject to foundational cryptographic assumptions about the uniqueness of hashes, may be comprehensive of an arbitrary number of transactions, is sufficient.
@@
Trusted computing is not necessary to protect users from loss or theft, either. To protect their assets, users may, at their option, choose devices that offer security features such as encryption, passcodes, or local biometrics to protect against theft, and they may make copies of their tokens to mitigate the risk of accidental loss of their devices. Users can retain copies of their tokens, and they can share copies of the tokens with the police if the devices containing them are lost, stolen, or compromised. Users can also upload copies of their tokens to non-custodial service providers, such as cloud storage services. As described in Section 4, double-spending is prevented by the requirement that every asset is bound to a specific integrity provider and a specific single-use key for specifying the next transaction on the asset.
@@
A much better way to address the limited set of cases in which “fully offline” payments (that is, those where neither the payer nor the payee have access to any suitable third parties or network infrastructure) are necessary would be to support the use of cash itself where cash continues to work well, such as in remote locations and environments with low connectivity, while avoiding all of the aforementioned problems associated with trusted hardware. To address the observed reduction of cash acceptance in urban environments [Mansuri and Vyas 2024, De Nederlandsche Bank 2025] by providing digital currency, whilst encouraging the provision of facilities for converting between cash and digital currency at the boundary of “unbuilt” environments, would be an entirely legitimate approach. In particular, if one potential benefit of CBDC in an age of waning cash infrastructure [^4] Evidence for the decline of cash is growing [Access to Cash Review 2019, Tisher 2020], and it is acknowledged as an important risk by central banks, including the ECB [European Commission 2023a]. is the ability to electronically transport money to remote population centres during crises [Goodell 2020], then there is no reason to require that central bank money distributed in this manner must remain digital. For example, service providers in a particular locality can provide physical tokens in exchange for CBDC, where the CBDC is transferred using one of the methods described earlier, much as a casino can provide chips in exchange for cash or bank money.
@@
None of this is to say that we should forbid limited groups of users from trusting certified devices, if that is what they freely choose, to conduct transactions without network connectivity to a third party. Such an approach is entirely compatible with an architecture that expects that the third party would generally operate as a network service. However, at a system level, requiring all users to purchase and carry certified devices, even if some users intend only to hold money offline so that they can transact it online, introduces costs and risks that are entirely avoidable. It is possible to build a common framework that supports both online and single-offline settlement involving assets held offline in free hardware devices and dual-offline settlement involving assets held in certified hardware devices. Oblivious management offers an opportunity to deliver on this promise by allowing both online service providers and certified hardware devices to serve as integrity providers for digital assets.
@@ 7 Conclusion
In the end, we must ask ourselves: What problem are we really trying to solve with digital currency? Manifestly, the problem is the declining use of cash. But this decline is really only present in environments with high network connectivity, wherein dual-offline payments are not necessary. For environments with low network connectivity, the decline of cash has not occurred, and cash continues to be the best solution for such environments.
@@
Many central banks, including the Bank of England, have recently affirmed a commitment to cash, while underscoring the risks associated with the decline of cash use in favour of online payment schemes such as card payment networks. This article has argued that the use cases that digital currency payments should be targeting are cases for which at least one counterparty is already online. To argue for digital currencies with the premise that support for dual-offline payments are its principal source of added value mainly supports the view that such products are solutions in search of a problem, as the UK House of Lords prominently suggested [House of Lords 2021].
@@
The purpose of digital wallets (and indeed, digital currency in general) is not primarily about enabling transactions in which both parties are offline. The purpose of digital wallets is to allow users to hold digital monetary bearer instruments directly, without a custodian. Put differently, this is a way for users to hold digital assets that they can exchange. By facilitating such payments, digital currency can protect the interests of consumers in all cases, including fully networked ones. If the counterparty has network connectivity (for example, an electronic point of sale device), then transactions involving digital wallets can be consummated without certified hardware. Even if we accept that fair exchange is impossible without the participation of a trusted third party, and that the trusted third party can participate via the network or via trusted hardware, it is important to underscore that a non-networked digital wallet can conduct a fair exchange with a networked counterparty, without the aid of special security mechanisms built into the wallet.
@@
It is also a mistake to assume the primary motivation for CBDC is to allow a government to phase out cash, or that custodial wallets are sufficient for all scenarios except the narrow case in which neither the sender nor the receiver have a communication line to a third party. Part of the reason that cash is successful is its role as a bearer instrument. Credit systems actually predated cash, and cash emerged as a partial substitute [Graeber 2011]. What the world does not yet have is an institutionally supported version of cash for the digital economy, and a digital currency aspiring to fill this role will require a non-custodial wallet architecture.
@@
Custodial wallets require network access as a rule, because consumers must be able to use their devices to access their custodial accounts. However, support for non-custodial wallets does not imply a design choice between network access and certified hardware in the general case. In particular, single-offline payments can make use of non-custodial wallets without certified hardware to hold money offline and transact it online, leveraging the network connectivity provided by POS devices presented to customers by merchants. To deny the salience of this use of wallets is to impede the development of a technology that can protect the privacy and autonomy of users while limiting the empowerment of big players in digital payments infrastructure. Ultimately, the digital wallets that support economic transactions should be seen not as access devices or proxies offered by custodians, but as tools freely chosen, and even built, by end-users to store and use digital money autonomously, on their own terms.
@@ Acknowledgements
The author thanks Professor Tomaso Aste for his continued support for his ongoing work on digital currencies and digital payment systems. The author thanks Christian Grothoff and Yuri Biondi for their generous and insightful feedback on earlier drafts of this article. The author also thanks his many colleagues in the international standards development community for a series of vigorous arguments that were instrumental to reaching the crux of this most contentious debate. The author also acknowledges the Future of Money Initiative at University College London, the Systemic Risk Centre at the London School of Economics, EPSRC, and the PETRAS Research Centre EP/S035362/1 for the FIRE Project.